Suppose the lengths of the pregnancies of a certain animal are approximately normally distributed with a mean of μ=188 days and a standard deviation of σ=13 days. What is the probability that a randomly selected pregnancy lasts less than 184 ​days?
Answer #1

Son. X= logw @ Prga X N( и, с*) = N(188, 13*) - 2 - Х - 4 X - 18 8- 15 No, 1) Рух< 18Ч) = Р(х - и 184-188 4 12 = P 24-. б. 3please.upvote